OpenStudy (anonymous):

how do i write an expression for P(X<=x) for the density function f9x)={x, 0<=x<=1

OpenStudy (ash2326):

\[P(X<=x)=\int_{-\infty}^{x} f(x) dx\]

OpenStudy (ash2326):

Here \[0\le x \le 1\] so \[P(X\le x)=\int_{0}^x f(x) dx\]

OpenStudy (ash2326):

@willlliam Do you get it?

OpenStudy (anonymous):

yes thank you very much, i really appreciate all the help.